Yegnaw Company produces and selis two types of yoga-trainıng products: how-to videotapes and a basic equipment set (blocks, strap, and small pillows). Last year, Yegnaw sold 14,500 videos and 7,250 equipment sets. Information on the two products is as follows: Video Set Equipment Price Br.17.40 Br.21.75 Variable cost per unit 5.80 8.70 Total fixed costs are Br.101,500. Required: Answer the following 1. What is the sales mix of videos and equipment sets 2. Compute weighted average contribution margin 3. Compute the break-even quantity of each product. 4. What is weighted average contribution margin ratio 5. What is the overall break-even sales revenue?

1
Expert's answer
2022-03-16T11:58:03-0400
  1. 14500+7250 = =21750
  2. (101500 - 17.50) = 101482.50
  3. 14500/3= 4833.33
  4. (101500 - 17.50) = 101482.50
  5. 14500+7250 = 21750